# Local Hybrid Retrieval-Augmented Document QA

### v1

@@ 2.2 Hybrid Retrieval Strategies
Dense retrieval using transformer-based embeddings excels at semantic similarity but struggles with out-of-vocabulary terms sentencetransformers; bgeembeddings. Sparse methods like BM25 provide precise lexical matching but miss conceptual relationships hybridretrieval. Recent work demonstrates that linear combination of sparse and dense scores often outperforms individual methods, though optimal weighting strategies remain empirically determined hybridretrieval.
@@ 3.3 Document Processing Pipeline
Documents are processed through specialized loaders supporting PDF (PyPDFLoader), CSV (CSVLoader), and JSON (JSONLoader) formats langchainjs. Content is chunked using RecursiveCharacterTextSplitter with optimized parameters (300-400 tokens, 30-50 token overlap) to maintain semantic coherence while enabling efficient retrieval. These values were selected empirically under constrained compute and time resources; a broader sensitivity sweep (e.g., 128–1024 tokens with learned dynamic merging) is deferred to future work once additional resources are available.
@@ 3.4 Local Language Model Integration
The generation pipeline constructs prompts that combine: (1) formatted chat history for conversational context, (2) retrieved document chunks as grounding context, and (3) the current user question. Ollama parameters are configured for balanced output quality (temperature: 0.6, top-p: 0.95, top-k: 40, max tokens: 1024), optimizing for factual accuracy while maintaining natural language fluency. This local inference approach provides sub-second generation latency on consumer hardware while maintaining strict data sovereignty requirements.
@@ 4.3 Hallucination Evaluation Protocol
We adopt an LLM-as-Judge framework to quantify hallucination and faithfulness. Note: While the production system operates 100% locally, we leverage Gemini’s API exclusively for offline evaluation to enable scalable assessment of 1,500+ query-answer pairs across three datasets; no production user data is transmitted externally. For each dataset, a stratified sample of 500 queries (uniform over question length tertiles) is selected. This size balances statistical precision and constrained evaluation budget: at observed rates (0.8%–6.2%) a Wilson 95% interval remains within roughly ±2 percentage points while keeping API cost and labeling time manageable under limited resources. Larger stratified expansions (e.g., 1k–2k) are deferred to future work when additional compute and budget are available. For every query we store: (i) user question, (ii) retrieved context chunks (top 5, concatenated with provenance IDs), and (iii) model answer under the optimal hybrid retriever.
@@ 5.5 Reliability Assessment
LLM-as-Judge evaluation using Gemini for both answer generation and hallucination detection across 500 queries per dataset llm_judge; hallucination_survey.
@@ 7 Conclusion
Key findings show that hybrid retrieval consistently outperforms single-method approaches, GPU acceleration yields major speedups, LLM-as-Judge reveals low error rates, and the modular design supports secure enterprise deployment.
The system demonstrates that effective RAG performance can be achieved without compromising data security, offering a practical solution for organizations requiring AI-powered document analysis while maintaining regulatory compliance and data sovereignty.
@@
Throughput is limited by synchronous retrieval and single-node design, causing tail latency under load. Future work includes async batching and sharded indices. Hallucination judgments rely on a single LLM-as-Judge; multi-judge ensembles and significance testing are planned. Benchmarks are English-only and web-skewed (SQuAD, MS MARCO, Natural Questions); multilingual and domain-specific evaluation is deferred. Security lacks RBAC, PII redaction, and retrieval poisoning detection—these will be addressed in future iterations through embedding-space anomaly scoring, signed chunk manifests, role-based access policies, and red-team audit telemetry.
@@
Although production inference is fully local, hallucination evaluation used Gemini’s API on benchmark-derived queries only—no user data was transmitted. This enables scalable assessment but introduces external dependency in evaluation, not deployment. Future work will explore local judge models (e.g., fine-tuned Llama) to eliminate this for organizations requiring fully offline pipelines.

### v2

@@ 2.2 Hybrid Retrieval Strategies
Dense retrieval using transformer-based embeddings excels at semantic similarity but struggles with out-of-vocabulary terms sentencetransformers; bgeembeddings. Sparse methods like BM25 provide precise lexical matching but miss conceptual relationships hybridretrieval. Recent work demonstrates that linear combination of sparse and dense scores often outperforms individual methods, though optimal weighting strategies remain empirically determined hybridretrieval. Our work applies established hybrid fusion techniques within a fully local deployment context, with the primary contribution being the integration and optimization of these methods for privacy-preserving enterprise environments rather than novel retrieval algorithms.
@@ 3.3 Document Processing Pipeline
Documents are processed through specialized loaders supporting PDF (PyPDFLoader), CSV (CSVLoader), and JSON (JSONLoader) formats langchainjs. Content is chunked using RecursiveCharacterTextSplitter with parameters (300-400 tokens, 30-50 token overlap) selected empirically under constrained compute and time resources. These values maintained semantic coherence while enabling efficient retrieval in preliminary tests; however, comprehensive sensitivity analysis across broader ranges (e.g., 128–1024 tokens with varying overlaps) represents a limitation of this work and is deferred to future investigation once additional resources are available.
@@ 3.4 Local Language Model Integration
The generation pipeline constructs prompts that combine: (1) formatted chat history for conversational context, (2) retrieved document chunks as grounding context, and (3) the current user question. Ollama parameters are configured for balanced output quality (temperature: 0.6, top-p: 0.95, top-k: 40, max tokens: 1024), optimizing for factual accuracy while maintaining natural language fluency. This local inference approach maintains strict data sovereignty requirements, with end-to-end query response times (retrieval + generation) averaging 2 seconds on consumer hardware.
@@ 4.3 Hallucination Evaluation Protocol
We adopt an LLM-as-Judge framework to quantify hallucination and faithfulness. Note: All answers in both production and evaluation are generated by Llama 3.2 via Ollama (fully local). Gemini’s API is used exclusively for offline hallucination judging (not generation) to enable scalable assessment of 1,500+ query-answer pairs across three datasets; no production user data is transmitted externally. For each dataset, a stratified sample of 500 queries (uniform over question length tertiles) is selected. This size balances statistical precision and constrained evaluation budget: at observed rates (0.8%–6.2%) a Wilson 95% interval remains within roughly ±2 percentage points while keeping API cost and labeling time manageable under limited resources. Larger stratified expansions (e.g., 1k–2k) are deferred to future work when additional compute and budget are available. For every query we store: (i) user question, (ii) retrieved context chunks (top 5, concatenated with provenance IDs), and (iii) model answer under the optimal hybrid retriever.
@@ 5.5 Reliability Assessment
LLM-as-Judge evaluation using Gemini for hallucination detection (answers generated by Llama 3.2) across 500 queries per dataset llm_judge; hallucination_survey.
@@ 7 Conclusion
Key findings show that hybrid retrieval consistently outperforms single-method approaches, GPU acceleration yields major speedups, LLM-as-Judge reveals low error rates, and the modular design supports secure enterprise deployment. While the hybrid fusion technique employs standard linear weighting rather than novel adaptive methods, the contribution lies in demonstrating that privacy-preserving local RAG can achieve competitive performance on established benchmarks.
The system demonstrates that effective RAG performance can be achieved without compromising data security during deployment, offering a practical solution for organizations requiring AI-powered document analysis while maintaining regulatory compliance and data sovereignty.
@@
Throughput is limited by synchronous retrieval and single-node design, causing tail latency under load. Future work includes async batching and sharded indices. Hallucination judgments rely on a single LLM-as-Judge without extensive human validation; multi-judge ensembles and human-annotated validation sets are planned. The hybrid strategy employs standard linear weighting without novel adaptive or query-specific fusion mechanisms; comparison with cross-encoder rerankers and dynamic hybrid methods (e.g., DAT, MoR) is deferred. Benchmarks are English-only and web-skewed (SQuAD, MS MARCO, Natural Questions); multilingual and domain-specific evaluation is needed. Chunk size sensitivity analysis (128–1024 tokens) was not exhaustively explored due to compute constraints. Security lacks RBAC, PII redaction, and retrieval poisoning detection—these will be addressed in future iterations through embedding-space anomaly scoring, signed chunk manifests, role-based access policies, and red-team audit telemetry.
@@
Although production inference is fully local (Llama 3.2 generates all answers via Ollama), hallucination evaluation used Gemini’s API for judging only—no user data was transmitted, and Gemini did not generate any answers. This creates an evaluation dependency but not a deployment dependency. The "fully local" claim applies strictly to the operational system; evaluation leverages external APIs for scalability. Future work will explore local judge models (e.g., fine-tuned Llama) and human validation to eliminate external evaluation dependencies for organizations requiring fully offline pipelines.